OpenStudy (anonymous):

Carissa and Louann both have savings accounts. Carissa has $250 in her account and deposits $80 per month. Louann has $1230 in her account and withdraws $60 per month. After how many months will they have equal amounts of money in their accounts? What amount of money will each have when it is the same amount? It will take months__. Then, they will both have $__ in their accounts. i am stumped on this one @iGreen

OpenStudy (igreen):

I answered this yesterday too..xD

OpenStudy (igreen):

We can create two expressions and solve them: Carissa: 250 + 80x Louann: 1230 - 60x Solve: 250 + 80x = 1230 - 60x Add 60x to both sides: 250 + 140x = 1230 Subtract 250 to both sides: 140x = 980 Can you divide 980 / 140? @NunChuckBoy3101

OpenStudy (anonymous):

yes it =7

OpenStudy (igreen):

Yep, so let's plug that back into one of the equations to find the second number: 250 + 80x 250 + 80(7) 250 + 560 Can you add those?

OpenStudy (anonymous):

250+80x=810......250+560=810

OpenStudy (igreen):

Yep..

OpenStudy (igreen):

So 7 goes in the first blank, and 810 goes in the second.

OpenStudy (igreen):

@NunChuckBoy3101

OpenStudy (anonymous):

ok thank you i wont bother you any more